OpenStudy (anonymous):

What is the area of sector BEC when CE = 9 yd?

OpenStudy (jdoe0001):

got pic?

OpenStudy (jdoe0001):

 \( \textit{area of a sector}=\cfrac{\theta \pi r^2}{360}\qquad \begin{cases} \theta\to angle\\ r\to radius \end{cases}\)

OpenStudy (anonymous):

Thanks!

OpenStudy (jdoe0001):

yw